# RoDUS: Robust Decomposition of Static and Dynamic Elements in Urban Scenes

### v1

@@ 3.2 Enabling Semantic Awareness
Figure 4: (b) The use of a trimmed kernel helps completely remove moving cars in the scene. (c, d, e) As the sky and road regions are badly reconstructed, applying class constraints can aid the learning process of the static field.
3.3 Loss Formulation
Robust loss. Our objective is to reduce the sensitivity of the static model to dynamic objects by incorporating supervision through a robust photometric loss. This loss can be formulated as an Iteratively Reweighted Least Squares (IRLS) problem:
where $\epsilon(\mathbf{r})=\|\hat{C}^{S}(\mathbf{r})-C(\mathbf{r})\|_{2}$ is the error residuals and $\omega(\epsilon)=\epsilon^{-1}\cdot\partial\kappa(\epsilon)/\partial\epsilon$ is the weight function of the residuals at the previous iteration. Among the family of kernels [1], we follow Sabour et al. [34] and choose a trimmed estimator [6] which classifies outliers based on a percentile threshold $\mathcal{T}_{\epsilon}$ calculated from the whole batch of rays. This choice ensures outliers are completely removed (i.e. hard weights). Furthermore, as every object has neighbor connections, they apply a spatial filter $\mathcal{F}$ on the residuals map to capture the smoothness, ensuring local support of the kernel:
@@ 3.3 Loss Formulation
3.4 Training strategy
The use of robust estimation plays an important role in rejecting inconsistency (dynamic objects in our case) and guiding the static branch without any prior knowledge. However, we observe that using a binary kernel alone with a fixed threshold, after reaching a certain point throughout the training, the robust loss does not contribute to the overall loss anymore. We refer to this phenomenon as “kernel saturation”, this causes several parts of the scene not to be learned completely. Adjusting the threshold $\mathcal{T}_{\epsilon}$ seems to be a trivial solution, but given the fact that outdoor scenes are more complicated to handle, there is always a trade-off between the number of distractors removed and the static scene quality. Particularly, high-frequency but static details, equivalently appear as large residuals are considered as outliers, while low-frequency but moving regions such as shadows (dark) which mostly share a similar color with the road (gray) are uncontrollably included as inliers. Figure 4 portrays this, as there are still parts of the scene (e.g. the truck on the right) that are not fully learned by the model.
In contrast, relying solely on reconstruction losses leads to several local optima during optimization due to the entanglement of geometry and appearance. We propose using the robust loss as an initialization step for the static model to learn the basic overall structures of the scene before delving into complete scene understanding. The initialization step guides the network to prioritize consistent (or confident) regions that further help stabilize the training process. Therefore, we start the model training with the robust loss (Eq. 5) and sky loss (Eq. 7) enabled for the first epoch, and subsequently enable the remaining losses. This serves as a foundation for disentangling the scene in later training stages.
@@ 3.4 Training strategy
Figure 5: We demonstrate RoDUS’s capabilities on multiple tasks. Notice that we successfully retain separation consistencies between all photometric, geometric, and semantic outputs. Legend: GT: Ground truth, $\hat{I}$: Composed RGB, $\hat{d}$: Composed depth, $I^{S}$: Static RGB, $\hat{d}^{S}$: Static depth, $\hat{I}^{D}$: Dynamic RGB (blended with a green background), $\hat{S}$: Composed segmentation map, $\hat{S}^{S}$: Static segmentation map, $M^{D}$: Motion segmentation.
@@ 4.1 Experimental Setup
[…]
@@
KITTI-360 RobustNeRF [${}^{{\dagger}}$ 19.50 0.722 0.282 - - -
[…]
@@
Pandaset RobustNeRF [${}^{{\dagger}}$ 23.41 0.764 0.197 - - -
[…]
@@
${}^{{\dagger}}$ As RobustNeRF rejects distractors during training, full image reconstruction metrics are not included.
GT
RobustNeRF [34]
N/AN/A
[…]
SUDS [43]
EmerNeRF [52]
RoDUS (ours)
Composed RGB
Static RGB
Dynamic RGB
[…]
@@ 4.3 Dynamic Objects Segmentation
[…]
GT image Refer to caption Refer to caption Refer to caption
@@
[…]
@@ 4.4 Ablation Study
Refer to caption w/o $\mathcal{L}_{road}$ Refer to caption w/o $\mathcal{L}_{robust}$ Refer to caption Full model static RGB Refer to caption Full model static depth Figure 8: Static depth map ablation
@@ 4.5 Limitation
Our approach can’t handle areas that lack observations or have been highly occluded most of the time within the sequence. This is particularly notable by large or slow-moving objects, illustrated in Fig. 9. Exploring a model that is capable of detecting and inpainting these missing regions based on prior knowledge will be interesting as a future research direction.
@@ 5 Conclusion
In this paper, we presented RoDUS, a method for disentangling dynamic scenes using neural representation. Our approach effectively separates moving components from the scene without the need for manual labeling. To enhance disentangling efficiency, we proposed a robust training strategy and incorporated semantic reasoning, resulting in high-quality background view synthesis and accurate segmentation. We conducted extensive experiments on challenging autonomous driving datasets, showcasing that RoDUS surpasses current state-of-the-art NeRF frameworks.
@@ 0.A.1 Data Processing
In our experiments, we downsample Pandaset images by a factor of 4 ($480\times 270$) and KITTI-360 by a factor of 2 ($704\times 188$). The selected scenes from KITTI-360 are summarized in Tab. 3. For Pandaset, we used sequences 023, 042, and 043. We specifically selected those scenes because they feature multiple dynamic objects while providing sufficient views of the background regions.
@@ 0.A.2 Architecture
We adopt a proposal-based coarse-to-fine sampling technique[2] using two proposal models with the number of samples being $[128,96]$ for the proposal models, and 48 for the final model. Our hash grids have the size of $2^{19}$ for $16$ levels with the base resolution and finest resolution of 16 and 2048 respectively. Additionally, our dynamic hash grid allows 4D inputs, which means that the intermediate feature vector is computed through a quadrilinear interpolation of 16 grid point vectors instead of the 8 used for the 3D hash grid. The static branch of our model is equipped with a 16-dimension per-frame embedding to adapt to varying lighting conditions.
@@ 0.A.4 Training
We trained RoDUS for 50 epochs using the Adam optimizer with a batch size of 200. The learning rate was set to $0.01$, and we used cosine decay to decrease it to $0.0001$. The rays are grouped in $15\times 15$ patches to suit DSSIM, planar regularization, and robust losses. Section 0.A.4 summarizes the hyperparameters used for the training of RoDUS, we do not require manual tuning for each scene, but we believe that such tuning could potentially further enhance performance.
Table 4: Training hyperparameters
@@
[…]
@@ Appendix 0.B Additional Results
Fig. 10 and Fig. 11 show additional qualitative comparisons between our RoDUS and previous methods in static-dynamic decomposition and scene reconstruction. Fig. 12 displays RoDUS’s performance on other different scenes, demonstrating that RoDUS’s dynamic branch is able to classify moving objects from parked vehicles or other static instances. Furthermore, RoDUS can be trained from noisy semantic annotations, leading to a scene-consistent semantic representation. A combination of multiple modalities also supports the model in reconstructing and segmenting thin structures and objects (e.g. light poles, and traffic signs).
GT
RobustNeRF [34]
N/AN/A
[…]
SUDS [43]
EmerNeRF [52]
Figure 10: Additional qualitative comparisons on Pandaset.
RoDUS (ours)
Composed RGB
Static RGB
Dynamic RGB
@@
RobustNeRF [34]
N/A
N/A
[…]
@@
GT image Refer to caption
RobustNeRF [34]
N/A
N/A
[…]
@@
Figure 12: Additional results on RoDUS’s capability.
@@ 0.C.1 Robust Kernel Threshold
We explain the choice of percentile threshold $\mathcal{T}_{\epsilon}$ used in RobustNeRF [34] and report qualitative results in Fig. 13. With $\mathcal{T}_{\epsilon}=0.7$, there are still a lot of structures that could not be learned (e.g. the far away building), when we increase $\mathcal{T}_{\epsilon}$ to $0.9$, the shadow regions started to appear. As a result, we choose $\mathcal{T}_{\epsilon}=0.75$ as it yields the best visual quality with the least artifacts. Correspondingly, the kernel saturates as we observed no improvement as the training continues. Notice that the model could not learn the road marking in any case.
@@
0.C.2 Shadow Handling

### v2

@@ 3.3 Loss Formulation
Figure 4: (b) Using a trimmed kernel effectively removes all moving cars from the scene. (c, d, e) As the sky and road regions are badly reconstructed, applying class constraints can aid the learning process of the static model.
@@
Figure 5: We demonstrate RoDUS’s capabilities on multiple tasks. Legend: GT: Ground truth, $\hat{I}$: Composed RGB, $\hat{d}$: Composed depth, $I^{S}$: Static RGB, $\hat{d}^{S}$: Static depth, $\hat{I}^{D}$: Dynamic RGB (blended with a green background), $\hat{S}$: Composed segmentation map, $\hat{S}^{S}$: Static segmentation map, $M^{D}$: Motion segmentation.
@@
3.4 Training Strategy
The use of robust estimation plays an important role in rejecting inconsistency (dynamic objects in our case) and guiding the static branch without any prior knowledge. However, we observe that when using a binary kernel alone with a fixed threshold, after reaching a certain point throughout the training, the robust loss does not contribute to the overall loss anymore. We refer to this phenomenon as “kernel saturation”, this causes several parts of the scene not to be learned completely. Adjusting the threshold $\mathcal{T}_{\epsilon}$ seems to be a trivial solution, but given the fact that outdoor scenes are more complicated to handle, there is always a trade-off between the number of distractors removed and the static scene quality. Particularly, high-frequency but static details, equivalently appear as large residuals are considered as outliers, while low-frequency but moving regions such as shadows (dark) which mostly share a similar color with the road (gray) are uncontrollably included as inliers. Fig. 6 portrays this, as there are still parts of the scene (e.g. road marking) that are not fully learned.
In contrast, relying solely on reconstruction losses leads to several local optima during optimization due to the entanglement of geometry and appearance. We propose using the robust loss as an initialization step for the static model to learn the basic overall structures of the scene before delving into complete scene understanding. The initialization step guides the network to prioritize consistent (or confident) regions that further help stabilize the training process. Therefore, we start the model training with the robust loss (Eq. 5) and sky loss (Eq. 7) enabled for the first epoch, and subsequently enable the remaining losses. By doing so, we can confidently select a low threshold without compromising the quality. This serves as a foundation for disentangling the scene in later training stages.
@@ 4.1 Experimental Setup
Baseline. We benchmark RoDUS against 4 different state-of-the-art methods for static scene extraction (RobustNeRF [34]) and dynamic scene rendering (D2NeRF [47], SUDS [43], and EmerNeRF [52]). Evaluations focus on two key factors: the quality of novel view synthesis of the static image and the segmentation masks of dynamic objects. While our primary focus is on static-dynamic disentanglement, RoDUS’s design allows it to represent dynamic scenes holistically so the quality of the overall image is also considered. We showcase the performance of our method on three representative scenarios in Fig. 5
@@
KITTI-360 RobustNeRF [34] - - - 19.50 0.722 0.282 - - -
D2NeRF [47] 23.86 0.688 0.232 18.70 0.588 0.309 12.04 4.67 7.41
@@
Pandaset RobustNeRF [34] - - - 23.41 0.764 0.197 - - -
D2NeRF [47] 25.25 0.692 0.291 23.38 0.682 0.292 - - -
@@
As RobustNeRF rejects distractors during training, full image reconstruction metrics are not included.
GT Refer to caption
RobustNeRF [34] N/A Refer to caption N/A
D2NeRF [47] Refer to caption Refer to caption Refer to caption
SUDS [43] Refer to caption Refer to caption Refer to caption
EmerNeRF [52] Refer to caption Refer to caption Refer to caption
RoDUS (ours) Refer to caption Refer to caption Refer to caption
Composed RGB Static RGB Dynamic RGB
Figure 6: Reconstruction and decomposition qualitative results comparisons on Pandaset. D2NeRF struggles in intricate driving contexts and produces poor results. SUDS faces difficulties in distinguishing between ego-motion and object-motion due to its reliance on optical flow supervision, leading to a large number of cloudy artifacts in the dynamic image to match the ground truth optical flows. RobustNeRF successfully reconstructs the background with no dynamic objects remaining, but it suffers from saturation, leading to degraded results (as explained in Sec. 3.3). We can see that the separation quality is still suboptimal. In contrast, RoDUS successfully decouples dynamic objects and achieves a clean background with no artifacts remaining.
@@ 4.3 Dynamic Objects Segmentation
For this task, we obtain the motion mask $M^{D}$ by rendering dynamic opacities and applying a threshold $\mathcal{T}_{O^{D}}=0.5$. We report mask Recall, IoU, and F1-score on training views. We only evaluate on KITTI-360 due to the absence of ground-truth instance masks in Pandaset. In Fig. 7, we explore various scenarios to understand how the ego-actor moves in relation to other moving objects and how it observes the scenes that could impact the decomposition. Particularly, case (a) is likely the easiest as the background is broadly observed for a long period, we can still notice a good separation in D2NeRF, SUDS’s results. In cases (b) and (c), when the ego actor is in motion, most methods suffer from incorrect segmentation, while our semantic guidance forces the model to learn the correct dynamic objects. SUDS tends to produce large dynamic areas, resulting in very high Recall but low IoU and F1, yet EmerNeRF exhibits sensitivity to lighting variations. Correspondingly, Tab. 1 indicates that RoDUS’s metrics outperform the others in both tasks.
GT Refer to caption Refer to caption Refer to caption
@@
D2NeRF [47] Refer to caption Refer to caption Refer to caption
@@ 4.4 Ablation Study
Table 2: Ablation study. Impact of our design choice on static NVS task.
@@
Full model 25.94 0.814 0.170
w/o $\mathcal{L}_{d}$0.80 -0.032 +0.023
w/o $\mathcal{L}_{sem}$1.67 -0.080 +0.078
w/o $\mathcal{L}_{robust}$0.91 -0.047 +0.067
w/o $\mathcal{L}_{road}$0.84 -0.035 +0.026
w/o $\mathcal{L}_{\sigma^{D}}$0.74 -0.030 +0.024
w/o FG mask -0.67 -0.035 +0.023
w/o FG mask
w/o $\mathcal{L}_{sem}$
w/o $\mathcal{L}_{robust}$
Full model
Figure 8: Components ablation on $\hat{I}^{S}$.
@@ 4.5 Limitation
In general, RoDUS works best with scenes where labels belonging to potentially movable objects are well-defined. Nonetheless, if the semantic ground truth is absent or heavily mislabeled, it moderately impacts performance. Moreover, our approach can’t handle areas that lack observations or have been highly occluded most of the time within the sequence. This is particularly notable by large or slow-moving objects in Fig. 9. Exploring a model that is capable of uncertainty modeling to incorrect or anomalous pixels and inpainting these missing regions based on prior knowledge will be interesting as a future research direction.
@@ 5 Conclusion
In this paper, we presented RoDUS, a method for disentangling dynamic scenes using neural representation. Our approach effectively separated moving components from the scene without the need for manual labeling. To enhance disentangling efficiency, we proposed a robust training strategy and incorporated semantic reasoning, resulting in high-quality background view synthesis and accurate segmentation. Extensive experiments on challenging autonomous driving datasets showcased that RoDUS surpasses current state-of-the-art NeRF frameworks.
@@ 0.A.1 Data Processing
In our experiments, we downsample Pandaset images by a factor of 4 ($480\times 270$) and KITTI-360 images by a factor of 2 ($704\times 188$). The selected scenes from KITTI-360 are summarized in Tab. 3. For Pandaset, we use sequences 023, 042, and 043. We specifically select those scenes because they feature multiple dynamic objects while providing sufficient views of the background regions.
@@ 0.A.2 Architecture
We adopt a proposal-based coarse-to-fine sampling technique[2] using two proposal models with the number of samples being $[128,96]$ for the proposal models, and 48 for the final model. Our hash grids have the size of $2^{19}$ for $16$ levels with the base resolution and finest resolution of 16 and 2048 respectively. Additionally, our dynamic hash grid allows 4D inputs, which means that the intermediate feature vector is computed through a quadrilinear interpolation of 16 grid point vectors instead of 8 used in the 3D hash grid. The static branch of our model is equipped with a 16-dimension per-frame embedding to adapt to varying lighting conditions.
@@ 0.A.4 Training
We train RoDUS for 50 epochs using the Adam optimizer with a batch size of 200. The learning rate is set to $0.01$, and decayed to $0.0001$. The rays are grouped in $15\times 15$ patches to suit DSSIM, planar regularization, and robust losses. Sec. 0.A.4 summarizes the hyperparameters used for the training of RoDUS, we do not require manual tuning for each scene, but we believe that such tuning could potentially further enhance performance.
Table 4: Training hyperparameters.
@@
Baseline Implementations. We adapt the code in the official repository of each method: RobustNeRF [^1] https://robustnerf.github.io/, D2NeRF [^2] https://github.com/ChikaYan/d2nerf, SUDS [^3] https://github.com/hturki/suds, EmerNeRF [^4] https://github.com/NVlabs/EmerNeRF and modify their dataloaders. We search for the best robust loss threshold $\mathcal{T}_{\epsilon}$ for RobustNeRF, while the rest are trained with the default parameters provided in the repositories.
@@ Appendix 0.B Additional Results
Fig. 10 and Fig. 11 show additional qualitative comparisons between our RoDUS and previous methods in static-dynamic decomposition and scene reconstruction. Fig. 12 displays RoDUS’s performance on other different scenes, demonstrating that RoDUS’s dynamic branch is able to classify moving objects from parked vehicles or other static instances. Furthermore, RoDUS can be trained from noisy semantic annotations, leading to a scene-consistent semantic representation. Our combination of multiple modalities also supports the model in reconstructing and segmenting thin structures and objects (e.g. light poles, and traffic signs).
@@
RobustNeRF [34] N/A Refer to caption N/A
D2NeRF [47] Refer to caption Refer to caption Refer to caption
@@
Composed RGB Static RGB Dynamic RGB
Figure 10: Additional qualitative comparisons on Pandaset.
GT Refer to caption
RobustNeRF [34] N/A Refer to caption N/A
D2NeRF [47] Refer to caption Refer to caption Refer to caption
SUDS [43] Refer to caption Refer to caption Refer to caption
EmerNeRF [52] Refer to caption Refer to caption Refer to caption
RoDUS (ours) Refer to caption Refer to caption Refer to caption
GT Refer to caption
RobustNeRF [34] N/A Refer to caption N/A
D2NeRF [47] Refer to caption Refer to caption Refer to caption
@@
Figure 12: Additional showcases of RoDUS’s capability.
@@ 0.C.1 Robust Kernel Threshold
We explain the choice of percentile threshold $\mathcal{T}_{\epsilon}$ used in RobustNeRF [34] and report qualitative results in Fig. 13. With $\mathcal{T}_{\epsilon}=0.7$, there are still a lot of structures that could not be learned (e.g. the far away building), when we increase $\mathcal{T}_{\epsilon}$ to $0.9$, the shadow regions start to appear. As a result, we select $\mathcal{T}_{\epsilon}=0.75$ for benchmarking as it yields the best visual quality with the least artifacts. Correspondingly, the kernel saturates and we observed no improvement as the training continues. Notice that the model still could not learn the road marking in any case.
@@
0.C.2 Various Challenging Scenarios
Fig. 14 illustrates RoDUS’s performance in low light (driving at night) and uneven illumination (driving into a shaded area) environments with high variations between views which are common in an autonomous driving context. It also highlights that the model could handle chaotic traffic and slow-moving entities (i.e. cyclists) as long as there is no significant occlusion.
Complex traffic Refer to caption Refer to caption Refer to caption
Cyclist Refer to caption Refer to caption Refer to caption
Low light Refer to caption Refer to caption Refer to caption
Uneven Illumination Refer to caption Refer to caption Refer to caption
Composed RGB Static RGB Dynamic RGB
Figure 14: RoDUS’s adaptivity on different scenarios.
0.C.3 Shadow Handling